Given that logcx = a and logcy = b:
Find in terms of a and b:
a) logc(c^2 x^4 y^3)
b) logc((c^3 x^2)/5rooty)
c) logy(x^2)


Answer by lwsshak3(11628)

Given that logcx = a and logcy = b:
Find in terms of a and b:
a) logc(c^2 x^4 y^3)
b) logc((c^3 x^2)/5rooty)
c) logy(x^2)
I will assume 5rooty=5√y
...
a) logc(c^2 x^4 y^3)
=2logc(c)+4logc(x)+3logc(y) (multiplication & power rule for logs)
=2(1)+4a+3b
=2+4a+3b
..
b) logc((c^3 x^2)/5√y)
=3logc(c)+2logc(x)-(logc(5)-1/2logc(y)
=3+2a-logc(5)-b/2